OpenStudy (anonymous):

If h(x) = 2x+1/5x+6, what is h(1/2)?

OpenStudy (anonymous):

everywhere u see an x in the equation, plug in 1/2

OpenStudy (anonymous):

what answer did you get???

OpenStudy (anonymous):

i want to see if i am correct :D

OpenStudy (anonymous):

2(1/2)+1/5(1/2)+6 1+7/10+6 77/10 7.7 that what u got?

OpenStudy (anonymous):

i got 1/17

OpenStudy (anonymous):

:(

OpenStudy (anonymous):

I'm rechecking my work...

OpenStudy (anonymous):

Is it (2x + 1) / (5x + 6) or 2x + (1/5x) + 6

OpenStudy (anonymous):

(2x+1)/(5x+6)

OpenStudy (anonymous):

ohhh! That's why

OpenStudy (anonymous):

0.23

OpenStudy (anonymous):

I get 4/17

OpenStudy (anonymous):

got2 go, helping someone w/ a problem

OpenStudy (anonymous):

4/17

OpenStudy (anonymous):

2(1/2) + 1 = 2 5(1/2) + 6 = 5/2 + 12/2 = 17/2 2/(17/2) = 2 * (2/17) = 4/17